integrate the following integral
∫ cos 3x cos 4x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int cos3xcos4xdx
I = \frac{1}{2}\int 2cos3xcos4xdx
I = \frac{1}{2}\int [cosx+cos7x]dx
I = \frac{1}{2}[\int cosxdx+\int cos7xdx]
I = \frac{1}{2}[sinx+\int cos7xdx]
I = \frac{1}{2}[sinx+\frac{sin7x}{7}]+c